### v1

@@ Abstract
Blockchain-based smart contracts offer a new take on credible commitment, where players can commit to actions in reaction to actions of others. Such reaction-function games extend on strategic games with players choosing reaction functions instead of strategies. We formalize a solution concept in terms of fixed points for such games, akin to Nash equilibrium, and prove equilibrium existence. Reaction functions can mimic “trigger” strategies from folk theorems on infinitely repeated games—but now in a one-shot setting.We introduce a refinement in terms of safe play. We apply our theoretical framework to symmetric investment games, which includes two prominent classes of games, namely weakest-link and public-good games. In both cases, we identify a safe and optimal reaction function. In this way, our findings highlight how blockchain-based commitment can overcome trust and free-riding barriers.
@@ 1 Introduction
Smart contracts are digital tools that allow agents to coordinate in a decentralized environment (Gans, 2021; Bakos and Halaburda, 2023; Gudmundsson et al., 2024). In short, a smart contract is a piece of computer code with a unique address on a blockchain. Agents with shared access to the blockchain can deploy such contracts, deposit funds to them, and hard-code how these funds are used in reaction to choices made by others. As such, an agent can let a contract act automatically on her behalf, in effect making contract design a strategic choice. Once deployed, a contract cannot be changed. In this way, smart contracts ensure commitment as strong as the cryptographic security of the network itself (e.g. Liang et al., 2024; Raskin, 2017).
We submit that smart contracts can be used as commitment devices when agents play a one-shot game using conditional strategies. In particular, the contract allows players to program reaction functions (i.e., actions chosen by the player contingent on the actions chosen by the other players), which in turn can be submitted to a coordinating contract that determines the outcome that is realized and executed. Such simple commitment devices are well suited to handle a broad class of symmetric investment games prominent in economics. To this point, we illustrate how reaction-function play allows players to overcome trust issues in weakest-link games and help players battle free-riding in public-good games.
As such, we follow up on a line of literature that mixes ideas from game theory and computer science, dating back to von Neumann and further developed through “metagames” (Howard, 1971), “program equilibrium” (Tennenholtz, 2004), and “commitment games” (Kalai et al., 2010). In short, the common element is that players’ strategies consists of commitment devices/programs which specify actions in response to the devices chosen by the others. In Tennenholtz (2004), players choose programs that specify a probability distribution over actions (mixed strategy) in response to every profile of programs chosen by the others. A program equilibrium is thus a Nash equilibrium (in pure strategies) in the game with strategy spaces consisting of programs rather than actions. The notion of program equilibrium allows a form of “folk theorem” in a one-shot context: any outcome with at least maxmin payoffs is achievable in program equilibrium (Theorem 2 in Tennenholtz, 2004). [^1] Further developments of program equilibrium include Monderer and Tennenholtz (2009), Oesterheld (2019), and Cooper et al. (2025). An analogous result is established in Kalai et al. (2010) using their notion of commitment games. In contrast to Tennenholtz (2004), their device specifies an action (pure strategy) in response to the devices chosen by others, and they allow for mixed-strategy equilibria in the commitment game.
Our reaction functions can be viewed as a simpler type of commitment device with the important difference that we condition on the actions (pure strategies) of the other players and not their devices. Like Tennenholtz (2004), we do not consider mixed strategies. While reaction functions are easier to formulate for the players and thereby more useful for practical implementation, they introduce the problem of determining the outcome of a profile of reaction functions as there can be zero, one, or multiple fixed points for a given profile of reaction functions. We introduce two assumptions to deal with this problem: players evaluate a profile by their preferred fixed point of the profile, and players prefer any outcome to a lack of fixed point (in which case we can imagine that players “try again”). Since, we still may have situations where players expect different fixed-points to be implemented on the basis of the same profile of reaction functions we further focus on unambiguous profiles for which there is a Pareto-superior fixed-point. With these assumptions, we define a reaction-function equilibrium as an unambiguous profile of reaction functions for which there is no beneficial unilateral deviation. In other words, no player can hope to realize a fixed point with higher payoff given the reaction functions of the other players.
In our first result, Theorem 3.1, we prove existence of reaction-function equilibrium for any reaction function-game. For two-player games, Theorem 3.2 characterizes the set of outcomes supported in reaction-function equilibrium as those where players get at least their maxmin payoff. This mimics the “folk theorems” obtained in Tennenholtz (2004) and Kalai et al. (2010). Yet, for games with more players, we can support essentially all outcomes in reaction-function equilibrium (Theorem 3.3). This is a consequence of that, with many players, reaction functions can be arranged such that fixed points are ruled out for any unilateral deviation. This calls for further refinements of the equilibrium concept. Loosely speaking, we then demonstrate that even though “bad” outcomes can be supported in artificial reactions, only “good” ones can be supported in compelling ways.
@@
A natural starting point are reaction functions that guarantee a player at least her maxmin payoff. In conventional strategic games, a player can always “play safe” by choosing the maxmin action. However, this may turn out to yield overly conservative outcomes as illustrated by the Prisoner’s dilemma game in which safe play coincides with the unique Pareto-dominated Nash equilibrium. Reaction functions allow us to do better in the sense that, if safe play constitutes a Nash equilibrium in the strategic game, then we can support any Pareto improvement in safe reaction-function equilibrium (Theorem 4.1). Yet, there are also games for which there does not exist a safe reaction-function equilibrium (Example 2).
To illustrate the potential of reaction functions as commitment devices, we zoom in on the class of symmetric investment games that includes the so-called weakest-link and public-good games. For these games, we first show that if all players’ reaction functions exhibit “conditional collaboration” (i.e., collaborate if the others do, and otherwise undercut the max investment of the others) we are able to support the Pareto-superior “high investment” outcome in reaction-function equilibrium (Theorem 5.2).
@@
In weakest-link games, each player wants to match the minimum investment of the others, and the higher this is, the better for all (e.g. Riedl et al., 2016). The Pareto superior outcome where all invest the highest amount is now a Nash equilibrium, as is every coordinated action where players invest the same amount, but the strategy to always play the lowest amount is “safe”. There appears to be an obvious solution—all invest the maximum amount—yet there is overwhelming experimental evidence that players fail to coordinate on the efficient equilibrium (Van Huyck et al., 1990; Battalio et al., 2001; Weber et al., 2001; Brandts and Cooper, 2006; Weber, 2006), with some rare exceptions (Engelmann and Normann, 2010; Riedl et al., 2016, 2021). The problem is that the mere belief or expectation that someone else will play the safe minimum investment action is all it takes to justify playing minimum investment yourself. In general, the more players involved, the harder it gets:even if player $1$ can credibly commit to the highest level, it will not sway player $2$ if there is doubt on player $3$’s intentions (e.g. Aumann, 1990). To overcome this, we consider the class of reaction functions where players exhibit ”conditional collaboration” and ”payoff consistency” (choose the same reaction across payoff equivalent action profiles) and single out the best-reply reaction function that matches the minimum investment of the others (Theorem 5.3). This reaction function is safe, and in the specific case of so-called ”high risk” weakest-link games (where investing higher amounts comes with greater risk) it is further singled out as being weakly dominant against monotone play (Theorem 5.4). This yields a simple, safe, and trustless way to obtain the Pareto-superior outcome and can overcome the inherent curse of mistrust.
In public-good games, each player’s investment benefits everybody, but is a net loss to the player herself (e.g. Bergstrom et al., 1986). Again, coordination around investing the maximal amount is welfare maximizing, but choosing to invest the minimum amount is safe and dominant. In Theorem 5.5 we show that ”conditional collaboration” and ”payoff consistency” this time singles out the reaction-function where a player reacts by investing the average (rounded down) of what the others have invested. In the particular case of high-risk public-good games, this reaction-function is further maximizing welfare subject to safe play (Theorem 5.6). This yields a simple and safe way to obtain the Pareto-superior outcome and overcome the inherent curse of free-riding.
In practice, setting up smart contracts in case of a weakest-link game, for example, will be rather simple: agents submit contracts (e.g. in the form of best-reply reaction functions that matches the lowest investment of the others) to a coordinating contract and deposits an amount of money corresponding to the maximum amount they prefer to invest. The coordinating contract takes the profile of deposits as the starting point and applies the reaction functions to find a fixed point. The fixed point (e.g. the profile corresponding to the smallest deposit among the players) is implemented in the sense that the contract withdraws the amount in question from each player’s deposit and returns potential surplus amounts. In such a setting, players can therefore safely deposit the amount corresponding to the greatest mutual benefit knowing that they will never collectively miscoordinate to their individual disadvantage.
@@
Reaction functions have been used extensively in experimental work since Selten (1967), there termed the “strategy method” (e.g., Selten et al., 1997). Eliciting reaction functions (“contribution tables” in Fischbacher et al., 2001; Thöni and Volk, 2018) yields a more comprehensive view of the experimental subject’s thought process and allows for instance testing “off-path” behavior (e.g. to compare lab data with subgame-perfection predictions). As it is demanding to map out a complete contingency plan, subjects are given plenty of time and may make more thoughtful decisions compared to the “direct-response method” (Brandts and Charness, 2011). In short, reaction functions are used to derive additional insights on subjects’ behavior in the strategic game being played. But the goal is merely to elicit reaction functions—there is no need to play them out against each other, especially with potential issues of non-existence or multiplicity of fixed points. For contrast, fixed points of the reaction-function profile are fundamental in our analysis. To illustrate the methodological difference, in Fischbacher et al. (2001) all subjects submit a reaction function and an unconditional strategy (corresponding to a constant reaction), after which one player is drawn at random and the realized outcome is uniquely determined (compare Remark 1) by that player’s reaction to the others’ unconditional strategies. An exception is Oechssler et al. (2022), who compute fixed points in a repeated public-good setting but limit to monotone step-function reactions (in at most two steps); see also Reischmann and Oechssler (2018).
@@
Outline. In Section 2, we introduce reaction-function games and compare with conventional strategic games. In Section 3, we focus on existence of reaction-function equilibrium and explore the set of outcomes that can be supported. In Section 4, we study the notion of “safe play” in reaction-function games. In Section 5, we present our main application to symmetric investment games, including weakest-link and public-good games. Section 6 closes with concluding remarks. All proofs are postponed to the Appendix.
@@ 2 Games and definitions
In this section, we introduce reaction-function games and relate them to conventional strategic games. In strategic games, formalized in Subsection 2.1, players individually select actions that jointly result in an outcome over which players have preferences. The classic solution concept is Nash equilibrium. In reaction-function games, players individually specify their reaction to the actions of other players and are able to commit to these reactions. Determining the outcome of a reaction-function profile is now more intricate and consequently defining the game and a natural equilibrium concept requires further assumptions. This is covered in Subsections 2.22.4. We assume players have complete information on the game.
@@ 2.1 Strategic games
There is a finite set of $n$ players $N$. Each player $i$ has a finite set of at least two actions $A_{i}$. An outcome $a\in A\equiv\times_{i}A_{i}$ is an action profile. Each player $i$’s preference is represented by a payoff function $u_{i}\colon A\to\mathbb{R}$, where $i$ derives payoff $u_{i}(a)$ at outcome $a$. The triple $(N,A,(u_{i})_{i})$ defines a strategic game. A Nash equilibrium is an action profile $a$ such that, for each player $i$ and action $a^{\prime}_{i}\in A_{i}$, $u_{i}(a)\geq u_{i}(a^{\prime}_{i},a_{-i})$. [^2] As usual, the subscript $-i$ projects onto the subspace relative to $N\setminus\{i\}$ and $a=(a_{i},a_{-i})$. Let $A_{-i}\equiv\times_{j\neq i}A_{j}$. Note that existence of a Nash equilibrium is not guaranteed as we do not consider mixed strategies.
@@ 2.2 Reaction functions
In reaction-function games, players choose reaction functions rather than actions. A reaction function $R_{i}\colon A_{-i}\to A_{i}$ specifies the action $R_{i}(a_{-i})\in A_{i}$ that player $i$ commits to play in reaction to the actions $a_{-i}\in A_{-i}$ of the other players. That is, we now map out a complete plan of what to do contingent on the actions of the others. Let $\mathcal{R}_{i}$ be player $i$’s reaction functions. As the number of actions is finite, $\mathcal{R}_{i}$ is also finite. A profile of reaction functions, a profile for short, is $R\in\mathcal{R}\equiv\times_{i\in N}\mathcal{R}_{i}$. For $a\in A$, let $R(a)\in A$ be the action profile in which each player $i$ plays action $R_{i}(a_{-i})$.
@@
The strategy space of reaction-function games is orders of magnitude larger than that for strategic games. With $n$ players, each with $m$ actions, there are $\left|{A}\right|=m^{n}$ outcomes (i.e., action profiles) whereas there are $\left|{\mathcal{R}}\right|=(m^{n})^{n}$ reaction-function profiles.
2.3 Executable outcomes in reaction-function games
Reaction functions are commitments to certain actions in reaction to the actions of the others. In this way, an outcome can be executed if and only if player commitments align: for each player $i$, the action $a_{i}$ must match $i$’s committed reaction to the actions $a_{-i}$ of the others. Put differently, outcome $a$ is a fixed point of the profile $R$ in the sense that, for each player $i$, $a_{i}=R_{i}(a_{-i})$. Let $\mathcal{E}(R)=\{a\in A:R(a)=a\}$ be the set of fixed points of $R$ (which may be empty). Our approach builds on the following two assumptions:

### v2

@@ Abstract
Blockchain-based smart contracts offer a new take on credible commitment, where players can commit to actions in reaction to actions of others. Such reaction-function games extend on strategic games with players choosing reaction functions instead of strategies. We formalize a solution concept in terms of fixed points for such games, akin to Nash equilibrium, and prove equilibrium existence. Reaction functions can mimic “trigger” strategies from folk theorems on infinitely repeated games—but now in a one-shot setting—for instance to support Pareto-improvements on Nash equilibrium outcomes. In some games, this can even be done through risk-free, safe reaction functions. We apply our theoretical framework to symmetric investment games, which includes two prominent classes of games, namely weakest-link and public-good games. In both cases, we highlight a particular safe and optimal reaction function. In this way, our findings highlight how blockchain-based commitment can help overcome trust and free-riding barriers.
@@ 1 Introduction
Smart contracts are digital tools that allow agents to coordinate in a decentralized environment (Gans, 2021; Bakos and Halaburda, 2023; Gudmundsson and Hougaard, 2024; Gudmundsson et al., 2024, 2025). In short, a smart contract is a piece of computer code with a unique address on a blockchain. Agents with shared access to the blockchain can deploy such contracts, deposit funds to them, and hard-code how these funds are used in reaction to choices made by others. As such, an agent can let a contract act automatically on her behalf, in effect making contract design a strategic choice. Once deployed, a contract cannot be changed. In this way, smart contracts ensure commitment as strong as the cryptographic security of the network itself (e.g. Liang et al., 2024; Raskin, 2017).
Smart contracts can thus serve as commitment devices in one-shot interactions: rather than choosing an action directly, a player can commit to a conditional action plan that prescribes what to do as a function of what others do. We model such plans as reaction functions (i.e., mappings from the actions chosen by the others to your own actions) and study the resulting reaction-function games. As such, we follow up on a line of literature that mixes ideas from game theory and computer science, dating back to von Neumann and further developed through “metagames” (Howard, 1971), “program equilibrium” (Tennenholtz, 2004), and “commitment games” (Kalai et al., 2010). Once players commit to reaction functions, an outcome is implementable when the commitments are mutually consistent: that is, the realized action profile is a fixed point of the reaction-function profile. A profile of reaction functions is a reaction-function equilibrium if no player can benefit by unilaterally changing her committed reaction function, taking into account the resulting fixed-point outcomes (the formal definition is found in Section 2).
Our main message is positive: we show that reaction-function play allows players to reach better outcomes. As smart-contract commitment is particularly viable in investment problems, made concrete through escrowed deposits and automated transfers, we devote particular attention to symmetric investment games. If investments are strategic complements (players want to invest more the more others invest), then reaction functions aid in solving coordination and trust issues; we illustrate this in weakest-link games (e.g. Riedl et al., 2016). On the other hand, if free-riding incentives are present (players want others to invest but prefer not to invest themselves), then reaction functions open for a form of conditional collaboration, this time illustrated in public-good games (e.g. Bergstrom et al., 1986).
In addition, we provide the following theoretical foundations:
  1. 1.
    We establish existence of equilibrium in every reaction-function game. This is possible without having to invoke, for instance, mixed strategies.
  2. 2.
    In two-player games, we characterize the outcomes that can be supported in reaction-function equilibrium. This provides a one-shot analogue of the familiar “folk theorem” logic. The key insight is that smart-contract commitment allows implementation of Pareto-efficient outcomes that are not Nash equilibria in the conventional strategic game.
  3. 3.
    With more than two players, the set of supportable outcomes becomes very large: reaction functions offer, in some sense, too much flexibility. In response, we develop a refinement based on safe play. In spirit, this rules out “implausible” commitments, much like equilibrium refinements rule out non-credible threats.
@@
In conventional strategic games, a player can always “play safe” by choosing a maxmin action that guarantees the highest possible worst-case payoff. The downside is that this may yield overly conservative outcomes. For instance, in the Prisoner’s dilemma, safe play leads to the unique Pareto-dominated Nash equilibrium. In reaction-function games, we can analogously restrict attention to safe reaction functions in the sense that, no matter the action profile of the others, the player commits to react in a way that ensures herself at least maxmin payoff. But reaction functions provide more flexibility: one can simultaneously play safe and support high-payoff outcomes. We find that, if safe play is a Nash equilibrium in the strategic game, then we can also support all potential Pareto improvements of this outcome in safe reaction-function equilibrium. For instance, Pareto-efficient cooperation can be made a safe reaction-function equilibrium in the Prisoner’s dilemma.
To emphasize the potential of smart-contract commitment, we hone in on symmetric investment games. These provide a natural ground for application of our theory. Smart contracts make conditional investment commitments concrete through escrowed deposits and automated transfers; “safe play” becomes simple and interpretable; and investment choices have a natural order (higher investment is “more”). Accordingly, focusing on monotone reaction functions both captures economically plausible behavior and makes fixed-point existence and computation straightforward. We leverage this structure and introduce two novel conditions on reaction functions—norm-proofness and payoff consistency—to obtain sharp, practical recommendations for two prominent classes of investment games.
@@
In weakest-link games, each player wants to match the minimum investment of the others, and the higher this is, the better for all (e.g. Riedl et al., 2016). There appears to be an obvious solution—all invest the maximum amount—yet there is overwhelming experimental evidence that players fail to coordinate on the efficient equilibrium (Van Huyck et al., 1990; Battalio et al., 2001; Weber et al., 2001; Brandts and Cooper, 2006; Weber, 2006), with some rare exceptions (Engelmann and Normann, 2010; Riedl et al., 2016, 2021). The mere belief or expectation that someone else will make the safe minimum investment is all it takes to justify making the minimum investment yourself (e.g. Aumann, 1990). Here, reaction-function commitment offers a straightforward resolution: each player commits to match the minimum investment of the others, which includes a promise of high investment if all others do the same. Interestingly, this type of “uniform commitment” has already found support elsewhere in the literature, for instance on climate negotiations (Weitzman, 2014; Schmidt and Ockenfels, 2021). With “high-risk” parameters, we can go yet further: the same reaction function is then singled out as weakly dominant against monotone play. Hence, reaction-function play provides a very compelling way to overcome the trust issues inherent in the weakest-link game.
In public-good games, each player’s investment benefits everybody but is a net loss to the player herself (e.g. Bergstrom et al., 1986; Falkinger et al., 2000; Croson, 2007). Again, coordination around investing the maximal amount is welfare maximizing, but choosing to invest the minimum amount is safe and now even dominant. Here, we identify reaction functions in which each player matches the average (rounded up or down) of the others’ investments. We argue further that rounding down is particularly appealing: it is safe for all parameter values and supports the high-investment outcome in a symmetric reaction-function equilibrium. Moreover, this type of monotone commitment makes free-riding self-defeating: if all others match the average rounded down, then a unilateral deviation to never invest collapses the profile to the zero-investment fixed point.
Implementation sketch. We briefly touch on how to take reaction-function theory to practice using smart-contract investment commitments. At a high level, each player deploys a contract detailing her reaction function and escrows a deposit that bounds her maximum investment. In this way, we combine commitment (the reaction function cannot be changed after deployment) with automated execution (escrowed funds can be transferred without further interaction). A simple coordinating contract can then
  1. 1.
    collect commitments and deposits,
  2. 2.
    compute a fixed point of the submitted reaction functions, and
  3. 3.
    execute the implied investments (returning unused deposits).
The main practical bottleneck is the fixed-point computation. However, for economic applications with natural classes of reaction functions (e.g., monotone functions on a finite lattice), simple iterative procedures will suffice and quickly pin down a fixed point.
Related literature. Our approach is related to a strand of literature in which players choose commitment devices or programs rather than actions, including metagames (Howard, 1971), program equilibrium (Tennenholtz, 2004; Monderer and Tennenholtz, 2009; Oesterheld, 2019; Cooper et al., 2025), and commitment games (Kalai et al., 2010). In these approaches, a player’s “commitment device” typically specifies what to do as a function of the devices chosen by others (see e.g. Kalai et al., 2010, Section 5, for details). As a consequence, even when the underlying action spaces are simple, the induced strategy space of devices is extremely rich. The objects to be chosen are very large—even in a two-player, two-action game, there will be an infinite number of devices/programs, and each device is itself made out of an infinite number of choices (what to do against every possible device). For contrast, our reaction functions condition plainly on actions of others; in a two-by-two game, there are only four reaction functions. In this way, strategic choices are kept parsimonious and interpretable, and practical implementation through simple contracts becomes tractable.
@@
Reaction functions are also superficially related to the experimental “strategy method” (Selten, 1967; Selten et al., 1997; Brandts and Charness, 2011), there sometimes known as “contribution tables” (Fischbacher et al., 2001; Thöni and Volk, 2018). However, in this field, the goal is merely to elicit reaction functions (e.g. to get a more comprehensive view of the experimental subject’s though process and test “off-path” behavior)—there is no need to play them out against each other. To illustrate the methodological difference, in Fischbacher et al. (2001) all subjects submit a reaction function and an unconditional strategy (corresponding to a constant reaction function), after which one player is drawn at random and the realized outcome is uniquely determined (compare Remark 1) by that player’s reaction to the others’ unconditional strategies. The method generates more data to analyze, but issues such as fixed-point computation, existence, and multiplicity are irrelevant. For contrast, these issues are at the core of our analysis. An exception is Oechssler et al. (2022), who develop a theoretical model and test its implications experimentally. They explore a repeated public-good game and limit to monotone two-step reaction functions (circumventing any fixed-point issues); see also Reischmann and Oechssler (2018).
Finally, there is a growing literature on smart contracts as commitment devices in dynamic games. Brzustowski et al. (2023) show that, giving a monopolist access to such credible dynamic contracts overturns the classic Coase conjecture and makes it possible to achieve payoff bounded away from the buyer’s lowest valuation for any choice of discount factor. Relatedly, so-called “Stackelberg games” have recently gained new attention in the computer science literature. Here, the follower can strategically deploy a reaction function that credibly commits to irrational responses in their subgame, thereby influencing the leader’s optimal choice of action to the follower’s own advantage (e.g. Hall-Andersen and Schwartzbach, 2021).
@@
Outline. In Section 2, we introduce reaction-function games and compare them to conventional strategic games. In Section 3, we explore the set of outcomes that can be supported in reaction-function equilibrium. In Section 4, we develop the notion of “safe” reaction functions. In Section 5, we present our main application to symmetric investment games. Section 6 closes with concluding remarks. Proofs are postponed to the Appendix.
@@ 2 Games and definitions
In this section, we introduce reaction-function games and relate them to conventional strategic games. In strategic games, formalized in Subsection 2.1, players individually select actions that jointly result in an outcome over which players have preferences. The classic solution concept is Nash equilibrium. In reaction-function games, players individually specify their reaction to the actions of other players and are able to commit to these reactions. Determining the outcome of a reaction-function profile is now more intricate and consequently defining the game and a natural equilibrium concept requires further elaboration. This is covered in Subsections 2.22.4. We assume players have complete information on the game.
@@ 2.1 Strategic games
There is a finite set of $n$ players $N$. Each player $i$ has a finite set of at least two actions $A_{i}$. An outcome $a\in A\equiv\times_{i}A_{i}$ is an action profile. Each player $i$’s preference is represented by a payoff function $u_{i}\colon A\to\mathbb{R}$, where $i$ derives payoff $u_{i}(a)$ at outcome $a$. The triple $(N,A,(u_{i})_{i})$ defines a strategic game. A Nash equilibrium is an action profile $a$ such that, for each player $i$ and action $a^{\prime}_{i}\in A_{i}$, $u_{i}(a)\geq u_{i}(a^{\prime}_{i},a_{-i})$. [^1] As usual, the subscript i-i projects onto the subspace relative to $N\setminus\{i\}$ and $a=(a_{i},a_{-i})$. Let $A_{-i}\equiv\times_{j\neq i}A_{j}$. Note that existence of a Nash equilibrium is not guaranteed as we do not consider mixed strategies.
@@ 2.2 Reaction functions
In reaction-function games, players choose reaction functions rather than actions. A reaction function $R_{i}\colon A_{-i}\to A_{i}$ specifies the action $R_{i}(a_{-i})\in A_{i}$ that player $i$ commits to play in reaction to the actions $a_{-i}\in A_{-i}$ of the other players. That is, players now map out a complete plan of what to do contingent on the actions of the others. Let $\mathcal{R}_{i}$ be the set of player $i$’s reaction functions. As action sets are finite, so is the set of reaction functions $\mathcal{R}_{i}$. A profile of reaction functions, a profile for short, is $R\in\mathcal{R}\equiv\times_{i\in N}\mathcal{R}_{i}$. For $a\in A$, let $R(a)\in A$ be the action profile in which each player $i$ plays action $R_{i}(a_{-i})$. Letting players choose reaction functions enlarges the strategy space considerably compared to strategic games: there are $\left|{\mathcal{R}}\right|=(m^{n})^{n}$ reaction-function profiles versus $\left|{A}\right|=m^{n}$ outcomes (i.e., action profiles).
@@
2.3 Fixed points in reaction-function games
Reaction functions are commitments to certain actions in reaction to the actions of the others. In this way, an outcome can be realized if and only if player commitments align: for each player $i$, the action $a_{i}$ must match $i$’s committed reaction to the actions $a_{-i}$ of the others. Put differently, outcome $a$ is a fixed point of the profile $R$ in the sense that, for each player $i$, $a_{i}=R_{i}(a_{-i})$. Let $\mathcal{E}(R)=\{a\in A:R(a)=a\}$ be the set of fixed points of $R$ (which may be empty). Our approach builds on the following two assumptions: